john rowed 12 miles down river took 2 hour but the return trip took 3 hours
how fast is john going with out the flow of water? and how fast is the current of water?

Found 2 solutions by mananth, graphmatics:
Answer by mananth(16949)

Boat speed =x mph
current speed =y mph
againstcurrent x-y 3 hours
with current x+y 2 hours

Distance = 12 miles distance= 12
t=d/r against current
12.00 / ( x - y )= 3.00

3.00 x - -3.00 y = 12.00 ....................1
downstream
12.00 / ( x + y )= 2.00
2.00 ( x + y ) = 12.00
2.00 x + 2.00 y = 12.00 ...............2
Multiply (1) by 2.00
Multiply (2) by 3.00
we get
6.00 x + -6.00 y = 24.00
6.00 x + 6.00 y = 36.00
12.00 x = 60.00
/ 12.00
x = 5 mph

plug value of x in (1)
3 x -3 y = 12
15 -3 y = 12
-3 y = 12 -15
-3 y = -3
y = 1 mph
Boat 5 mph
current 1 mph

Answer by graphmatics(170)

Let J be John's speed and let S be the stream's speed.
12/(J+S)=2
12/(J-S)=3

12=2J+2S
12=3J-3S

36=6J+6S
-24=-6J+6S

12=12S
S=1
12=2J+2(1)
J=5
So John's speed is 5mph and the current speed is 1mph.